could some one please help me with this problem.
A bus leaves a station at 1 p.m.;taveling west at an average rate of 44mi/h.One hour later a second bus leaves the same station,traveling east at a rate of 48mi/h.At what time will the two buses be 274 mi apart?

Answer by maxchuie(2)

you can use the formula {d=rt}
44x+48(x-1)=274
44x+48x-48=274
9x-48+48=274+48
92x=322
x=3 1/2
1:00pm +3 1/2 hrs = 4:30pm